OpenStudy (anonymous):

( 4√5) (3√6) my answer 4√30 any other way to write this?

OpenStudy (anonymous):

I don't think you've considered all the factors. Your radical part is correct, but you left out the factor of three in the second factor group. Keeping that in mind,\[( 4\sqrt5) (3\sqrt6) =4*3\sqrt5\sqrt6=12\sqrt{30}\]

OpenStudy (anonymous):

The factorization of 30 indicates that radical term won't simplify further.

OpenStudy (anonymous):

understood thank you

OpenStudy (anonymous):

You're welcome.